OpenStudy (laddiusmaximus):

how is cos(pi/4)= 1/sqrt2?? I look on the unit circle and its sqrt2/2 Ive been doing this for a few hours so maybe my brain needs a break.

OpenStudy (anonymous):

its the same thing

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{2} }\]rationalize it by multiplying by root 2 on denominator and nominator

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{2} } *\frac{ \sqrt{2} }{ \sqrt{2} }\]

OpenStudy (anonymous):

they do that because it's not nice to have roots in the denominator.